# An Analysis of the Impact of Gold Open Access Publications in Computer Science

### v1

@@ 6.1 Network Centrality Analysis
Estimating centrality in this publisher network is complicated because the volume of publications differs across publishers. In the period in question (2015 onwards), the dataset contains  250k Elsevier papers. For comparison, there are 120k MDPI papers and 37k from ACM. So the high-volume publishers will dominate centrality measures based on raw citation counts. Therefore, we consider the ratio of in-links to out-links for our analysis. This is shown in the heatmap in Figure 5. We are interested in citations into a node so the numbers indicate the ratio of citations from the column publisher to the row publisher. Larger numbers (green) in the column for a publisher indicate a higher influence. For example, the bottom left cell indicates that on average an IEEE paper receives 6.01 as many citations from a Hindawi paper as a Hindawi paper received from an IEEE paper. Likewise the disparity between an ACM and MDPI is 3.29; ACM papers receive 3.29 as many links from MDPI as MDPI papers receive from ACM. And, in general, the larger values shown in the lower left quadrant indicate that traditional publishers receive many more citations from Gold OA publishers than Gold OA publishers receive from traditional publishers.
@@
Given that Figure 5 represents a weighted network of the normalised citation flows between publishers, we can quantify the centrality of nodes in this network using eigenvector centrality. These scores are shown in Figure 6. While eigenvector centrality scores are presented here, the Katz centrality scores are similar. This effectively sums up the picture in Figure 5 and shows a clear difference between the traditional publishers (IEEE, ACM, Elsevier, Springer), which have eigenvector centrality scores between 0.2 and over 0.7, compared to the much lower centrality scores (<0.2) of the Gold OA publishers. Despite the publication volume of MDPI and IEEE Access they do not have a significant impact. It also highlights that IEEE and IEEE Access are very different.
@@ 6.2 Citations and Influential Citations
In our post-2015 dataset, there are 850,368 papers with 17,121,561 citations. Of these, 824,002 are marked as influential citations, that is 4.8% of the total. This is lower than the 14.6% figure in the original work by Valenzuela and Etzioni. There are several possible explanations for this difference. It might be due to the original work covering a different domain (computational linguistics). Or it could be because it is based on pre-2015 papers with more influential citations. Or their training data may be biased towards papers with at least some influential citations; as an aside, in our dataset, if we only consider papers with at least 1 influential citation, then approximately 12% of citations are influential, which more closely matches the Valenzuela and Etzioni number. Figure 7 shows how this fraction varies across publishers. At the high end, ACM papers are above 6%, and MDPI is at the bottom with 3.5%. Three of the bottom four publishers are Gold OA publishers.
@@
In Figure 8 we look at the status of the citing papers (not the papers receiving the cites) and, in particular, their fraction of influential cites. In this case, the variation is less, but the overall trend remains the same. At the top, papers that cite ACM papers tend to have higher fractions of influential citations compared with papers that cite MDPI or Hindawi papers, for example. Once again, three of the bottom four publishers by this measure are Gold OA publishers.
Looking at overall citation counts in Figure 8, we see that the IEEE papers come out on top with an average of almost 30 citations per paper. Elsevier and the ACM are placed second and third, with IEEE Access in fourth position. Indeed, IEEE Access fares better than the other Gold OA publishers in this citation analysis regarding overall citation count and the proportion of influential cites.
@@ 6.2.1 Statistical Significance
Given the volume of data behind the results presented in Figures 7 and 8 it would be expected that the differences we see are statistically significant. We have tested the statistical significance of the differences in mean citation and mean influential citation counts for the eight publishers. First, we run one-way ANOVA tests to check if any pair of means is different. For both ANOVAs, the $p$-values are effectively zero suggesting that at least some of the publishers have different population means. This allows us to run a Tukey honestly significant difference (HSD) test to make pairwise comparisons between publishers. The detailed results are presented in an Appendix to the repository version of this paper. Some headline results are as follows:
@@ 7 Conclusions
The CS publication landscape has changed radically in recent years. Journal output has increased from 20,000 to 120,000 papers per year from 2000 to 2023. This is partly due to a push to improve publication availability via OA publication models and, today, a significant proportion of all papers are from Gold Open Access publishers. However, there are concerns about a Gold OA downside, because the economics of OA publishing may incentivise quantity over quality. In this paper, we find evidence to support this:
  • Gold OA publishers have experienced much higher growth rates than their traditional counterparts.
  • On average Gold OA papers in CS receive fewer citations than they produce; in network centrality terms Gold OA publishers have low status.
  • Gold OA papers receive less influential citations than traditional journals. The citations they receive are from papers that have fewer influential citations than average.
@@
In summary, given the increased volume and visibility of Gold OA papers it may be important to educate researchers, especially young, early-career researchers—those most acutely exposed to the pressure of publish-or-perish—about the significant changes in the publication landscape. Moreover, the evidence reported here justifies changes in how publication quality is assessed to properly account for these landscape changes by considering authorship and impact more carefully.

### v2

@@ 6.1 Network Centrality Analysis
Estimating centrality in this publisher network is complicated because the volume of publications differs across publishers. In the period in question (2015 onwards), the dataset contains  250k Elsevier papers. For comparison, there are 120k MDPI papers and 37k from ACM. So the high-volume publishers will dominate centrality measures based on raw citation counts. Therefore, we consider the ratio of in-links to out-links for our analysis. This is shown in the heatmap in Figure 5. We are interested in citations into a node so the numbers indicate the ratio of citations from the column publisher to the row publisher. Larger numbers (green) in the column for a publisher indicate a higher influence. For example, the bottom left cell indicates that on average an IEEE paper receives 6.01 as many citations from a Hindawi paper as a Hindawi paper received from an IEEE paper. Likewise the disparity between an ACM and MDPI is 3.29; ACM papers receive 3.29 as many links from MDPI as MDPI papers receive from ACM. And, in general, the larger values shown in the lower left quadrant indicate that traditional publishers receive many more citations from New OA publishers than New OA publishers receive from traditional publishers.
@@
Given that Figure 5 represents a weighted network of the normalised citation flows between publishers, we can quantify the centrality of nodes in this network using eigenvector centrality. These scores are shown in Figure 6. While eigenvector centrality scores are presented here, the Katz centrality scores are similar. This effectively sums up the picture in Figure 5 and shows a clear difference between the traditional publishers (IEEE, ACM, Elsevier, Springer), which have eigenvector centrality scores between 0.2 and over 0.7, compared to the much lower centrality scores (<0.2) of the New OA publishers. Despite the publication volume of MDPI and IEEE Access they do not have a significant impact. It also highlights that IEEE and IEEE Access are very different.
@@ 6.2 Citations and Influential Citations
In our post-2015 dataset, there are 850,368 papers with 17,121,561 citations. Of these, 824,002 are marked as influential citations, that is 4.8% of the total. This is lower than the 14.6% figure in the original work by Valenzuela and Etzioni. There are several possible explanations for this difference. It might be due to the original work covering a different domain (computational linguistics). Or it could be because it is based on pre-2015 papers with more influential citations. Or their training data may be biased towards papers with at least some influential citations; as an aside, in our dataset, if we only consider papers with at least 1 influential citation, then approximately 12% of citations are influential, which more closely matches the Valenzuela and Etzioni number. Figure 7 shows how this fraction varies across publishers. At the high end, ACM papers are above 6%, and MDPI is at the bottom with 3.5%. Three of the bottom four publishers are New OA publishers.
@@
In Figure 8 we look at the status of the citing papers (not the papers receiving the cites) and, in particular, their fraction of influential cites. In this case, the variation is less, but the overall trend remains the same. At the top, papers that cite ACM papers tend to have higher fractions of influential citations compared with papers that cite MDPI or Hindawi papers, for example. Once again, three of the bottom four publishers by this measure are New OA publishers.
Looking at overall citation counts in Figure 8, we see that the IEEE papers come out on top with an average of almost 30 citations per paper. Elsevier and the ACM are placed second and third, with IEEE Access in fourth position. Indeed, IEEE Access fares better than the other New OA publishers in this citation analysis regarding overall citation count and the proportion of influential cites.
@@ 6.2.1 Statistical Significance
Given the volume of data behind the results presented in Figures 7 and 8 it would be expected that the differences we see are statistically significant. We have tested the statistical significance of the differences in mean citation and mean influential citation counts for the eight publishers. First, we run one-way ANOVA tests to check if any pair of means is different. For both ANOVAs, the $p$-values are effectively zero suggesting that at least some of the publishers have different population means. This allows us to run a Tukey honestly significant difference (HSD) test to make pairwise comparisons between publishers. The detailed results are presented in the Appendix. Some headline results are as follows:
@@
7 More Robust Metrics
It is clear from our analysis that the validity of established publication metrics such as the $h$-index and FWCI is open to question because they are insensitive to author count or the source of a citation. We have seen that the average author count has doubled since 2000; there is now a strong case for taking author count into consideration. The eigenvector centrality and influential citation analyses in section 6 show that it is not reasonable to treat all citations the same as happens with the $h$-index and FWCI.
Based on the review of citation metrics presented in section 3 we already have mechanisms for addressing these issues. There is well-established research on how to normalise metrics such as the $h$-index to adjust for author count [6]. A recent paper in PLOS ONE[11] demonstrates that an author’s $h$-index is no longer a good indicator of research quality as evidenced by awards that indicate recognition by the scientific community. Instead a fractional allocation of $h$-index based on author count is a better indicator.
If citations are to be used as an indicator of quality then it should be clear from the analysis here that not all citations have the same status. Metrics based on network centrality such as the Eigenfactor Score and the Author Influence Score [5] do scale citations based on influence and provide a better indicator of the impact of a research article. As an example, when the SJR score mentioned in section 3.3 is used to compare MDPI Sensors with the Communications of the ACM it reveals an interesting picture: according to Scopus, both journals have similar cites per paper, yet the SJR scores are respectively 0.79 and 2.96. Thus, this SJR score captures the differences illustrated in Figures 5 and 6.
It is worth noting that the Clarivate Journal Citation Reports do include the Eigenfactor Score and Author Influence scores that help to quantify the status of citations. However, at present these are provided as secondary resources; their wider use would address the issues highlighted here.
@@ 8 Conclusions
The CS publication landscape has changed radically in recent years. Journal output has increased from 20,000 to 120,000 papers per year from 2000 to 2023. This is partly due to a push to improve publication availability via OA publication models, and today, a significant proportion of all papers are from New OA publishers. However, there are concerns about a New OA downside, because the economics of OA publishing may incentivise quantity over quality. In this paper, we find evidence to support this:
  • New OA publishers have experienced much higher growth rates than their traditional counterparts.
  • On average New OA papers in CS receive fewer citations than they produce; in network centrality terms New OA publishers have low status.
  • New OA papers receive less influential citations than traditional journals. The citations they receive are from papers that have fewer influential citations than average.
  • OA papers from traditional publishers do not suffer in this way, in fact there is evidence that they attract more citations than their non-OA counterparts.
@@
In summary, given the increased volume and visibility of New OA papers, it may be important to educate researchers, especially young, early-career researchers—those most acutely exposed to the pressure of publish-or-perish—about the significant changes in the publication landscape. Moreover, the evidence reported here justifies changes in how publication quality is assessed to properly account for these landscape changes by considering authorship and impact more carefully.